OpenStudy (anonymous.):

Anyone good with financial algebra ? I attempted to solve the question but my answer is not matching none of my answer choices :(

OpenStudy (anonymous.):

Kendra’s income puts her in the top tax bracket (35%) last year. During the same year, she earned $760 in dividends and $490 in coupons on a corporate bond. Three years ago, she purchased a common stock for $835. She sold the same stock last year for $1180. What is the total Kendra will pay in taxes on last year’s investments? $337.25 $223.25 $406.25 $165.75

OpenStudy (anonymous.):

1180-490 = 690 * 0.35 = 241.5 ???????

OpenStudy (anonymous):

you have the equation right...i would check your math tho...maybe you did somthing wrong and you didn't notice...if you've already done that, i can try to solve it

OpenStudy (anonymous):

if I had to say one...i would put B as ur answer

OpenStudy (anonymous.):

Thank you @turtleluver

OpenStudy (anonymous):

it was 337. 25